OpenStudy (anonymous):

Solve. 2x2 + x = 15.

OpenStudy (anonymous):

To solve this, we should first set right hand side of the eqt to be 0 2x^2 + x -15= 0 (2x-5)(x+3)=0 2x-5 =0 or x+3 =0 x=5/2 or -3

OpenStudy (anonymous):

-3